# FLTrojan: Privacy Leakage Attacks against Federated Language Models through Selective Weight Tampering

### v3

@@ 4.1 Threat model
In our settings, both the adversary and the victim(s) are members of the client group participating in FL to fine-tune the LLM. We assume that the adversary has some context of the victim’s training data, specifically, the non-sensitive prefixes of the privacy-sensitive texts they want to extract from the victim’s local dataset. This assumption is reasonable in practice because texts containing sensitive information (phone numbers, login credentials, etc.) often have some predictable formats, e.g., ‘For any queries, reach out to the manager at xxx-xxx-xxxx’, ‘Verify your credit card number details: xxxx-xxxx-xxxx-xxxx.’ However, for the sake of a more inclusive threat model, we also perform our attacks with inaccurate prefixes, i.e., prefixes with perturbations. Apart from that, we consider two independent cases of the adversary’s capability: (1) the adversary is a stand-alone malicious client in the FL setting who does not collude with the server and observe the global model in each round, and (2) optionally, the malicious client can collaborate with the server which enables it to exploit a specific victim’s local model updates. Without server cooperation, the adversary extracts certain types of sensitive data without targeting a specific client, and subsequently, all the users having such types of data in their local datasets fall victim to this attack. In contrast, with server cooperation, they can also target a particular victim and compromise their privacy. However, unlike many existing works that consider the server as an active adversary [51, 19, 18], we assume a more practical attack surface where the server is passive and helps the malicious participants by only sharing some meta information. This is a more practical setting since the server has some accountability, and it could cause a loss of its reputation if it runs active attacks during the FL training. Table 1 shows a high-level taxonomy of our attack methods. We delve into all the mentioned attack attributes in the later parts of this section. Figure 4 presents a high-level overview of our attack methods. There are two vital steps in our general attack pipeline: (1) Victim round identification (VRI): The adversary identifies the FL training rounds where victims with privacy-sensitive text of the adversary’s interest participated and (2) Maximizing data memorization (MDM): After the FL training is complete, the adversary manipulates the identified victim rounds’ models to maximize the memorization of sensitive data.
4.2 Attacks Without Server Cooperation
4.2.1 Victim Round Identification (VRI)
@@
As mentioned earlier, the attacker aims to extract some specific sensitive texts that follow certain patterns from the victims’ local datasets. If multiple clients’ local updates involve such private data of the attacker’s interest, all of them fall victim to the attack. The objective of the VRI step is to identify all the training rounds in which such victims participate. As mentioned in Section 1, LLMs tend to forget specific samples from an individual client’s training data as they tend to generalize the overall data distribution during the training process. To better realize this phenomenon, we exploit the exposure metric from [8] (definition details in Appendix D) and compute the average exposure (EXP) of the global model snapshots with regard to some victims’ private data for all rounds of FL. As shown in Figure 2(a), the value of EXP is higher for the rounds where the victim(s) participated (i.e., victim rounds, marked with blue dots) compared to the other rounds where they did not. Furthermore, EXP gradually decreases between two victim rounds, and more obviously, there is a persistent fall in the value of EXP after round 650 since the server did not select any victim in any round afterward (or any victim may not have participated). Based on our observation of this significant difference between the EXP values in victim and non-victim rounds, we want to deduce the following hypothesis:
@@
Step 1: The adversary stores the global model snapshots, i.e., $G_{i}$ for $i={1,\ldots,N}$, where N is the number of training rounds.
@@
We similarly calculate the norm of weight differences for $L_{\star}$ between $G$ and each $G_{i}$ and denote it as $\delta_{i}$. Our goal is to identify whether $G_{i}$ is a victim round where $i={1,\ldots,N}$.
@@
Step 4: Based on hypothesis 2, our key intuition of this final step is that if $G_{i}$ is a victim round snapshot, $\delta_{s}$ and $\delta_{i}$ would have a similar pattern in their distribution due to the presence of sensitive sequences in both of their training data, whereas $\delta_{r}$ and $\delta_{i}$ would have a significantly different distribution. For a better understanding of this phenomenon, refer to Figures 1(a) and 1(b). Hence, in this step, we perform two pairwise t-tests between $\delta_{i}$ and $\delta_{r}$, and between $\delta_{i}$ and $\delta_{s}$, respectively. For a predefined significance level, if the test with $\delta_{r}$ rejects the null hypothesis ($H_{0}$) and the test with $\delta_{s}$ accepts it, we mark $G_{i}$ as a victim snapshot. Otherwise, $G_{i}$ is categorized as a non-victim snapshot. To verify that such an approach works, we found the T-test between $\delta_{s}$ and $\|\Delta W_{\star}\|$ of any victim model almost always (95% cases) derive a p-value greater than the significance level, failing to reject the null hypothesis, which suggests an equivalent relationship between these two models. A detailed evaluation of our VRI method is provided in Section 5.4.1 It is important to note that $\mathcal{D}_{sen}$ has no overlap with clients’ private data. They are just some dummy out-of-distribution texts that have a similar pattern as the sensitive texts. More details on the curation of $\mathcal{D}_{sen}$ and $\mathcal{D}_{reg}$ are provided in the Attack Settings of Section 5.1.2. We formally present our VRI technique in Algorithm 1.
Algorithm 1 Victim Round Identification (VRI)
2:Output: Whether $G_{i}$ is a $\widehat{G_{i}}$ instance or not.
3:Fine-tune $G$ with $\mathcal{D}_{reg}$ and obtain $G_{r}$
4:Fine-tune $G$ with $\mathcal{D}_{reg}\cup D_{sen}$ and obtain $G_{s}$
5:$\delta_{r}\leftarrow(\|W(f_{G_{r}}(\mathcal{D}_{reg}))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
6:$\delta_{s}\leftarrow(\|W(f_{G_{s}}(\mathcal{D}_{reg}\cup\mathcal{D}_{sen}))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
7:$\delta_{i}\leftarrow(\|W(f_{G_{i}}(\cdot))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
8:$t_{r}:t$-$test(\delta_{i},\delta_{r},\alpha)$
9:$t_{s}:t$-$test(\delta_{i},\delta_{s},\alpha)$
10:if $t_{r}$ rejects $H_{0}$ AND $t_{s}$ accepts $H_{0}$ then
11:    $G_{i}$ is a $\widehat{G_{i}}$ instance
12:else
13:    $G_{i}$ is not a $\widehat{G_{i}}$ instance
14:end if
@@
4.2.2 Maximizing Data Memorization
After we identify victim snapshots ($\widehat{G_{i}}$), the next step is to maximize the victim models’ memorization of the private data based on hypothesis 1 by transforming their weights in certain layers ($W_{\star}$) as shown in Equation 1. We design two approaches to achieve this:
@@ 4.2.2 Maximizing Data Memorization
Maximizing data memorization with Selective Weights Optimization (SWO): We saw in the VRI step, for a victim round global snapshot $\widehat{G_{i}}$, the t-test between $\delta_{i}$ and $\delta_{r}$ rejects the null hypothesis. Hence we hypothesize that there must be a distinction in $W_{\star}$ between $\widehat{G_{i}}$ and $G_{r}$, which we denote as $\delta_{v}$.
@@
The optimization process can be performed using any conventional optimizer (e.g., SGD, Adam). The selective weights $W_{1}\in W^{\widehat{G_{i}}}_{\star}$ are updated iteratively over a predefined number of steps. The optimizer is reset at each step, and the new difference ($\Delta W_{new}$) as well as the cost ($\mathcal{L}$) are calculated with the updated $W_{1}$. We take the negative cost measure in line 8 since we aim to maximize it. Additionally, the optimization may include backpropagation with gradient clipping [53] and early stopping [55] to prevent exploding gradients [53] and running extra steps while updating the weights. In Figure 2(b) and 2(c), we show the contrast before and after applying SWO on a victim snapshot ($\widehat{G_{i}}$), respectively. It is evident that the norm of $\Delta W_{\star}$ has increased. Also Figure 3(a) indicates that the perplexity scores of the sensitive texts significantly decrease after applying SWO, whereas the scores remain almost as before in the case of plain English texts. Hence, we can deduce that with the selectively optimized weights, $\widehat{G_{i}}$ becomes more vulnerable to privacy leakage of the sensitive sequences without hurting its utility. Experimental results in Section 5 also advocate the above phenomenon.
@@
16:$\mathcal{M}\leftarrow\textsc{Learn\_Transform}(G,\mathcal{D}_{reg},\mathcal{D}_{sen},n)$
@@
Algorithm 3 gives a simplified demonstration of this transformation learning technique. Firstly, we need to generate $n$ (number of samples to train the supervised model $\mathcal{M}$) copies of $G_{r}$ and $G_{s}$ by fine-tuning $G$ with different sets of $\mathcal{D}_{reg}$ and $\mathcal{D}_{sen}$, respectively. Here, $D_{sen}$ does not need to be from the same distribution as the victim’s private data because according to hypothesis 1, the weights in $L_{\star}$ undergo similar changes for any out-of-distribution data. We get $n$ samples of $W_{\star}$ from both $G_{r}$ and $G_{s}$, denoted as $W_{r}$ and $W_{s}$, respectively. Finally, we train the supervised model $\mathcal{M}$ with $W_{r}$ as input and $W_{s}$ as output. Once $\mathcal{M}$ is trained, we can feed $W_{\star}$ of $\widehat{G_{i}}$ to $\mathcal{M}$ and get the updated weights. In Figure 2(b) and 2(d), We show the contrast before and after applying WTL on a victim snapshot ($\widehat{G_{i}}$) respectively. Besides increasing the magnitude of impact for $L_{\star}$ compared to Figure 2(c), it also shows a better alignment with the trend shown in Figure 13(b). Besides, in Figure 3(b), we observe that the language model’s memorization of the sensitive texts increases (perplexity goes down) by a higher margin with WTL when compared with SWO in Figure 3(a) and the model’s utility is also preserved.
4.3 Attacks With Server Cooperation
@@
The server can be honest-but-curious, a threat model that has been widely investigated in prior works [80, 70, 3]. In such cases, the server can passively cooperate with the adversary to enhance the victim participants’ privacy leakage but does not actively run the attack to remain stealthy. The advantage of including the server under the adversary is twofold. Firstly, it is important to note that the victim/non-victim round information is readily available to the parameter server. Although we have designed a victim-round identification mechanism as described earlier in Section 4.2.1, it may incur some errors in practice (see Table 2). In contrast, the server can provide the adversary with ground-truth victim round information– thus making the privacy leakage attacks more successful. Secondly, in the FL setting, the adversary receives global model updates from the parameter server in every iteration. As discussed in Section 4.2.2, the adversary aims to maximize the memorization of the victim’s privacy-sensitive data in those global model updates. However, it can be easily inferred without loss of generality that the impact of the victim’s data will be higher on its local model’s weights compared to the global model’s weights because local updates of all participants go through federated averaging, which weakens the victim’s impact on the global model. In other words, direct access to the local models may provide a greater advantage to the adversary in leaking the victim’s privacy-sensitive information. Therefore, the server can share the victim’s local model updates with the malicious clients, who can then maximize its memorization of private data using the techniques described in Section 4.2.2.
4.4 Static and Dynamic Modes of Attacks
@@
4.4.1 Static Mode
In this mode, the adversary does not actively interfere with the FL training process but instead passively exploits the model after the fine-tuning is completed. It involves two actions:
@@ 4.4.1 Static Mode
❏ After the entire training is complete, the adversary attempts to maximize the memorization of all identified victim snapshots using one of the MDM techniques discussed in Section 4.2.2.
4.4.2 Dynamic Mode
In this mode, the adversary maliciously alters its local model updates to exacerbate the memorization of the victims’ private data during the FL training. Its tasks involve the following two actions:
@@ 4.4.2 Dynamic Mode
Figure 4: Overview of our attack flow. First, we identify the training rounds the victim participated. If the server does not cooperate with the adversary, we use the Victim Round Identification algorithm. Next, we update the victim models to maximize the memorization of sensitive information using Maximizing Data Memorization algorithms. Finally, we prompt the model with crafted prefixes to retrieve sensitive information. More details about the algorithms can be found in Section 4.
4.5 Evaluating Privacy Leakage
4.5.1 Data Reconstruction
@@
After the VRI and MDM steps, our data reconstruction attack extracts private data from the victim models by prompting with the associated context using beam search [73]. Since our primary objective is to generate text with zero diversity and maximum resemblance to the training data, greedy decoding may seem to be a preferred approach. However, experiments show greedy decoding is subject to trivial memorization [9], e.g., repeated tokens and substrings, which often do not result in useful generations. Our experiments have shown that using the beam search method with a small beam size (2-5) is more effective in generating private texts within the training data. This particular generation scheme causes a significantly higher leakage of private texts than the others. We take top-3 accuracy, i.e., the three highest-scoring outputs produced by the beam search for each reconstruction are considered as the candidate data points leaked from the training data (i.e., total 200 canaries*3=600 reconstructions) for our evaluation and some sample reconstructions are provided in Table 6.
4.5.2 Membership Inference
@@
Besides data reconstruction, we also aim to evaluate whether our proposed strategies enhance the performance of membership inference (MI) [62] attacks. For this purpose, we adopt the reference-based attack proposed in [48], which involves calculating the percentage of private training samples that are correctly classified as training members out of a pool of both training and validation samples. To determine if a sample was part of the training set, two models are used: a target model, which is one of the victim models in our case, and a reference model. By comparing the likelihood probabilities from both models, a ratio is calculated, and if this ratio is below a certain threshold, the sample is considered part of the training set. We put the implementation details in the Appendix (Section G.1) due to space constraints.
5 Evaluation
5.1 Experiment Setup
5.1.1 Dataset Description

### v4

@@ IV-A Threat model
Without server cooperation, the adversary extracts certain types of sensitive data without targeting a specific client, and consequently, all the users having such types of data in their local datasets fall victim to this attack. In contrast, with server cooperation, they can also target a particular victim and compromise their privacy. However, unlike many existing works that consider the server as an active adversary [nguyen2023active, fowl2022decepticons, fowl2021robbing], we assume a more practical attack surface where the server is passive and helps the malicious participants by only sharing some meta information. This is a more practical setting since the server has some accountability, and it could cause a loss of its reputation if it runs active attacks during the FL training. Table I shows a high-level taxonomy of our attack methods. We delve into all the mentioned attack attributes in the later parts of this section. Figure 4 presents a high-level overview of our attack methods. There are two vital steps in our general attack pipeline: (1) Victim round identification (VRI): The adversary identifies the FL training rounds where victims with privacy-sensitive text of the adversary’s interest participated and (2) Maximizing data memorization (MDM): After the FL training is complete, the adversary manipulates the identified victim rounds’ models to maximize the memorization of sensitive data.
IV-B Victim Round Identification (VRI)
@@
As mentioned earlier, the attacker aims to extract some specific sensitive texts that follow certain patterns from the victims’ local datasets. If multiple clients’ local updates involve such private data of the attacker’s interest, all of them fall victim to the attack. The objective of the VRI step is to identify some of the training rounds (not necessarily all) in which such victims participate. As mentioned in Section I, LLMs tend to forget specific samples from an individual client’s training data as they tend to generalize the overall data distribution during the training process. To better realize this phenomenon, we exploit the exposure metric from [carlini2019secret] (definition details in Appendix D) and compute the average exposure (EXP) of the global model snapshots with regard to some victims’ private data for all rounds of FL. As shown in Figure 2(a), the value of EXP is higher for the rounds where the victim(s) participated (i.e., victim rounds, marked with blue dots) compared to the other rounds where they did not. Furthermore, EXP gradually decreases between two victim rounds, and more obviously, there is a persistent fall in the value of EXP after round 650 since the server did not select any victim in any round afterward (or any victim may not have participated). Based on our observation of this significant difference between the EXP values in victim and non-victim rounds, we want to deduce the following hypothesis:
@@
Step 1: The adversary stores the global model snapshots, $G_{i}$, where $i$ represents any of the $N$ training rounds when the attacker participates.
@@
We similarly calculate the norm of weight differences for $L_{\star}$ between $G$ and each $G_{i}$ and denote it as $\delta_{i}$. Our goal is to identify whether $G_{i}$ is a victim round.
@@
Step 4: Based on hypothesis 2, our key intuition of this final step is that if $G_{i}$ is a victim round snapshot, $\delta_{s}$ and $\delta_{i}$ would have a similar pattern in their distribution due to the presence of sensitive sequences in both of their training data, whereas $\delta_{r}$ and $\delta_{i}$ would have a significantly different distribution. For a better understanding, refer to Figures 1(a) and 1(b). Hence, in this step, we perform two pairwise t-tests between $\delta_{i}$ and $\delta_{r}$, and between $\delta_{i}$ and $\delta_{s}$, respectively. For a predefined significance level, if the test with $\delta_{r}$ rejects the null hypothesis ($H_{0}$) and the test with $\delta_{s}$ accepts it, we mark $G_{i}$ as a victim snapshot. Otherwise, $G_{i}$ is categorized as a non-victim snapshot. To verify that such an approach works, we found the T-test between $\delta_{s}$ and $\|\Delta W_{\star}\|$ of any victim model almost always (95% cases) derive a p-value greater than the significance level, failing to reject the null hypothesis, which suggests an equivalent relationship between these two models. A detailed evaluation of our VRI method is provided in Section VI-A1. It is important to note that $\mathcal{D}_{sen}$ has no overlap with clients’ private data. They are just some dummy out-of-distribution texts that have a similar pattern as the sensitive texts. More details on the curation of $\mathcal{D}_{sen}$ and $\mathcal{D}_{reg}$ are provided in the Attack Settings of Section V-B. We formally present our VRI technique in the Appendix (Algorithm 3).
@@
IV-C Maximizing Data Memorization
After we identify victim snapshots ($\widehat{G_{i}}$), the next step is to maximize the victim models’ memorization of the private data based on hypothesis 1 by transforming their selected weights ($W_{\star}$) as shown in Equation 1. We design two approaches to achieve this:
@@ IV-C Maximizing Data Memorization
Maximizing data memorization with Selective Weights Optimization (SWO): We saw in the VRI step, for a victim round global snapshot $\widehat{G_{i}}$, the t-test between $\delta_{i}$ and $\delta_{r}$ rejects the null hypothesis. Hence we hypothesize that there must be a distinction in $W_{\star}$ between $\widehat{G_{i}}$ and $G_{r}$, denoted as $\delta_{v}$.
@@
The optimization process can be performed using any conventional optimizer (e.g., SGD, Adam). The selective weights $W_{1}\in W^{\widehat{G_{i}}}_{\star}$ are updated iteratively over a predefined number of steps. The optimizer is reset at each step, and the new difference ($\Delta W_{new}$) as well as the cost ($\mathcal{L}$) are calculated with the updated $W_{1}$. We take the negative cost measure in line 8 since we aim to maximize it. Additionally, the optimization may include backpropagation with gradient clipping [pascanu2013difficulty] and early stopping [prechelt2002early] to prevent exploding gradients [pascanu2013difficulty] and running extra steps during the update. In Figure 2(b) and 2(c), we show the contrast before and after applying SWO on a victim snapshot ($\widehat{G_{i}}$), respectively. It is evident that the norm of $\Delta W_{\star}$ has increased. Also Figure 3(a) indicates that the perplexity scores of the sensitive texts significantly decrease after applying SWO, whereas the scores remain almost as before in the case of plain English texts. Hence, we can deduce that with the selectively optimized weights, $\widehat{G_{i}}$ becomes more vulnerable to privacy leakage of the sensitive sequences without hurting its utility. Experimental results in Section VI also advocate the above phenomenon.
@@
[…]
@@
Algorithm 2 gives a simplified demonstration of this transformation learning technique. Firstly, we need to generate $n$ (number of samples to train the supervised model $\mathcal{M}$) copies of $G_{r}$ and $G_{s}$ by fine-tuning $G$ with different sets of $\mathcal{D}_{reg}$ and $\mathcal{D}_{sen}$, respectively. Here, $D_{sen}$ does not need to be from the same distribution as the victim’s private data because according to hypothesis 1, the weights in $L_{\star}$ undergo similar changes for any out-of-distribution data. We get $n$ samples of $W_{\star}$ from both $G_{r}$ and $G_{s}$, denoted as $W_{r}$ and $W_{s}$, respectively. Finally, we train the supervised model $\mathcal{M}$ with $W_{r}$ as input and $W_{s}$ as output. Once $\mathcal{M}$ is trained, we can feed $W_{\star}$ of $\widehat{G_{i}}$ to $\mathcal{M}$ and get the updated weights. In Figure 2(b) and 2(d), We show the contrast before and after applying WTL on a victim snapshot ($\widehat{G_{i}}$). Besides increasing the magnitude of impact for $L_{\star}$ compared to Figure 2(c), it also shows a better alignment with the trend shown in Figure 2(b). Besides, in Figure 3(b), we observe that the language model’s memorization of the sensitive texts increases (perplexity goes down) by a higher margin with WTL when compared with SWO in Figure 3(a) and the model’s utility is also preserved.
IV-D Optional: Attacks With Server Collusion
@@
The server can be honest-but-curious, a threat model that has been widely investigated in prior works [zhao2020idlg, wang_hbc_server, lamp]. In such cases, the server can passively cooperate with the adversary to enhance the victim participants’ privacy leakage but does not actively run the attack to remain stealthy. The advantage of including the server under the adversary is twofold. Firstly, it is important to note that the victim/non-victim round information is readily available to the parameter server. Although we have designed a victim-round identification mechanism as described earlier in Section IV-B, it may incur some errors in practice (see Table II). In contrast, the server can provide the adversary with ground-truth victim round information– thus making the privacy leakage attacks more successful. Secondly, in the FL setting, the adversary receives global model updates from the parameter server in every iteration. As discussed in Section IV-C, the adversary aims to maximize the memorization of the victim’s privacy-sensitive data in those global model updates. However, it can be easily inferred without loss of generality that the impact of the victim’s data will be higher on its local model’s weights compared to the global model’s weights because local updates of all participants go through federated averaging, which weakens the victim’s impact on the global model. In other words, direct access to the local models may provide a greater advantage to the adversary in leaking the victim’s privacy-sensitive information. Therefore, the server can share the victim’s local model updates with the malicious clients, who can then maximize its memorization of private data using the techniques described in Section IV-C.
IV-E Static and Dynamic Modes of Attacks
@@
IV-E1 Static Mode
In this mode, the adversary does not actively interfere during the FL training but instead passively exploits the model after the fine-tuning is completed. It involves two actions:
@@ IV-E1 Static Mode
❏After the entire training is complete, the adversary attempts to maximize the memorization of all identified victim snapshots using one of the MDM techniques discussed in Section IV-C.
IV-E2 Dynamic Mode
In this mode, the adversary maliciously alters its local model updates to magnify the memorization of the victims’ private data during the FL training. Its tasks involve the following two actions:
@@ IV-E2 Dynamic Mode
Figure 4: Overview of our attack flow. First, we identify the training rounds the victim participated. If the server does not cooperate with the adversary, we use the Victim Round Identification algorithm. Next, we update the victim models to maximize the memorization of sensitive information using Maximizing Data Memorization algorithms. Finally, we prompt the model with crafted prefixes to retrieve sensitive information. More details about the algorithms can be found in Section IV.
IV-F Evaluating Privacy Leakage
IV-F1 Data Reconstruction
@@
After the VRI and MDM steps, our data reconstruction attack extracts private data from the victim models by prompting with the associated context using beam search [wiseman2016sequence]. Since our primary objective is to generate text with zero diversity and maximum resemblance to the training data, greedy decoding may seem to be a preferred approach. However, experiments show greedy decoding is subject to trivial memorization [carlini2021extracting], e.g., repeated tokens and substrings, which often do not result in useful generations. Our experiments have shown that using the beam search method with a small beam size (2-5) is more effective in generating private texts within the training data. This particular generation scheme causes a significantly higher leakage of private texts than the others. We take top-3 accuracy, i.e., the three highest-scoring outputs produced by the beam search for each reconstruction are considered as the candidate data points leaked from the training data (i.e., total 200 canaries*3=600 reconstructions) for our evaluation and some sample reconstructions are provided in Table VI.
IV-F2 Membership Inference
@@
Besides data reconstruction, we also aim to evaluate whether our proposed strategies enhance the performance of membership inference (MI) [shokri2017membership] attacks. For this purpose, we adopt the reference-based attack proposed in [mireshghallah2022quantifying], which involves calculating the percentage of private training samples that are correctly classified as training members out of a pool of both training and validation samples. To determine if a sample was part of the training set, two models are used: a target model, which is one of the victim models in our case, and a reference model. By comparing the likelihood probabilities from both models, a ratio is calculated, and if this ratio is below a certain threshold, the sample is considered part of the training set. We put the implementation details in the Appendix (Section G-A) due to space constraints.
V Experiment Setup
V-A Dataset Description